# Inference-Time Scaling of Discrete Diffusion Models via Importance Weighting and Optimal Proposal Design

### v3

@@ D.2 Additional Experimental Results
D.2.2 Additional Results on Language Modelling
@@
D.2.3 Additional Results on Biology Design
@@
D.2.4 Additional Results on Image Generation
Tables˜10 and 10 demonstrate additional results on enhancing CFG with the proposed SMC methods. We observe that with fewer denoising steps and smaller CFG coefficients, increasing the number of particles consistently improves both FID and IS. In contrast, when using more denoising steps and larger CFG coefficients, adding particles leads to higher IS but worse FID. This behavior aligns with our expectations. Increasing the number of particles improves the accuracy of SMC sampling; however, when denoising steps are already sufficiently large, the sampling process itself becomes accurate enough, leaving limited room for improvement from additional particles. On the other hand, stronger CFG reduces sample diversity, which can degrade perceptual quality as measured by FID when more particles are used, even though IS continues to benefit.
We further compare different sampling schemes (see Section˜C.1 for details) in Tables˜10, 10 and 10. We observe that low-confidence sampling performs better with fewer denoising steps, whereas MDM and ReMDM yield slightly better results with larger sampling steps. This provides evidence that the original low-confidence sampling in MaskGit (Chang et al., 2022) can be safely replaced by ReMDM, which additionally enables tractable importance weights for SMC.
@@ D.2.4 Additional Results on Image Generation
D.2.5 More Qualitative Results with Generated Samples
@@
  • In Figure˜12, we visualise the generated samples using different methods on ImageReward.
  • In Figure˜13, we visualise the generated samples on HPSv2.
  • In Figure˜14, we visualise the generated samples on Aesthetic Score.
  • In Figure˜15, we visualise the generated samples on ImageReward.
  • In Figure˜16, we demonstrate the generated toxic text using different methods.

### v4

@@ D.2 Additional Experimental Results
Figure 12: Normalised effective sample size (ESS) and reward of the particles across timesteps using different proposals without resampling.
Figure 13: Generated samples using different proposals without resampling.
D.2.2 ESS and Reward Traces for Different Proposals
Monitoring the effective sample size (ESS) provides a useful diagnostic of particle diversity over the SMC algorithm. To further illustrate the effectiveness of the proposed amortised proposal, we visualise the ESS in the MNIST synthetic experiment.
In this experiment, we evaluate three different proposals: $\text{Prop}_{\text{pretrained}}$, $\text{Prop}_{\text{grad}}$, and $\text{Prop}_{\text{amot}}$, corresponding to the pretrained diffusion proposal, the first-order approximated proposal, and the learned amortised proposal, respectively. Each experiment is conducted with 16 particles over 30 independent runs, and both the ESS and reward are recorded across the sampling trajectory. Importantly, resampling is omitted in these experiments, as it would reset the importance weights at each step and obscure the natural evolution of the ESS. Omitting resampling allows ESS to serve as a clearer measure of each proposal’s intrinsic ability to maintain particle diversity.
The ESS and reward trajectories are shown in Figure˜13, alongside the generated samples illustrated in Figure˜13. The results indicate that $\text{Prop}_{\text{amot}}$ consistently achieves the highest ESS and reward during sampling, demonstrating the effectiveness of the log-variance minimisation objective in learning an approximately optimal proposal. In contrast, while $\text{Prop}_{\text{grad}}$ achieves higher reward values than $\text{Prop}_{\text{pretrained}}$, it exhibits lower ESS, which is expected given that the first-order approximated proposal introduces bias relative to the optimal proposal and is therefore more prone to reward hacking. Regarding the generated samples, $\text{Prop}_{\text{grad}}$ fails to produce high-quality images, whereas $\text{Prop}_{\text{amot}}$ consistently generates visually coherent and realistic samples. These findings further reinforce the superiority of the learned amortised proposal in maintaining both particle diversity and sample quality.
D.2.3 Additional Results on Language Modelling
@@
We further compare our methods to SVDD (Li et al., 2024), which performs importance sampling at every unmasked step while aggressively maintaining only a single particle and using a pretrained diffusion model as its proposal distribution. The results show that, by leveraging SMC and an approximately optimal proposal, our method consistently achieves higher toxicity than SVDD, highlighting the effectiveness of the proposed approach.
@@
D.2.4 Additional Results on Biology Design
@@
We also compare with two additional baselines, SGDD (Chu et al., 2025) and SVDD (Li et al., 2024), to better contextualise the behaviour of our SMC methods. It is noteworthy that SGDD is a sampler restricted to uniform noising processes, while other methods in Table˜5 use masked diffusions. Although SGDD attains higher predicted activity and ATAC accuracy, its substantially weaker performance on correlation indicates mode-collapse behaviour, suggesting that it overfits to a narrow region of sequence space and struggles to generate diverse samples. SVDD shows a complementary pattern: while it attains strong predicted activity, its lower ATAC accuracy points to reward-hacking tendencies. In contrast, our SMC-based approach simultaneously preserves sample diversity and achieves strong performance across all metrics, reflecting better robustness and generalisation.
D.2.5 Additional Results on Image Generation
Tables˜11 and 11 demonstrate additional results on enhancing CFG with the proposed SMC methods. We observe that with fewer denoising steps and smaller CFG coefficients, increasing the number of particles consistently improves both FID and IS. In contrast, when using more denoising steps and larger CFG coefficients, adding particles leads to higher IS but worse FID. This behavior aligns with our expectations. Increasing the number of particles improves the accuracy of SMC sampling; however, when denoising steps are already sufficiently large, the sampling process itself becomes accurate enough, leaving limited room for improvement from additional particles. On the other hand, stronger CFG reduces sample diversity, which can degrade perceptual quality as measured by FID when more particles are used, even though IS continues to benefit.
We further compare different sampling schemes (see Section˜C.1 for details) in Tables˜11, 11 and 11. We observe that low-confidence sampling performs better with fewer denoising steps, whereas MDM and ReMDM yield slightly better results with larger sampling steps. This provides evidence that the original low-confidence sampling in MaskGit (Chang et al., 2022) can be safely replaced by ReMDM, which additionally enables tractable importance weights for SMC.
@@ D.2.5 Additional Results on Image Generation
SGDD ($\beta$ = 30) 8.85(0.07) 90.9(0.00) 0.470(0.014) 0.466(0.015) -263(1.6)
SGDD ($\beta$ = 50) 9.32(0.04) 96.4(0.01) 0.370(0.010) 0.398(0.001) -269(0.1)
SVDD (N = 8) 6.57(0.01) 67.4(0.01) 0.813(0.009) 0.753(0.011) -258(0.2)
SVDD (N = 16) 6.89(0.04) 84.3(0.01) 0.891(0.009) 0.834(0.011) -260(0.2)
@@
(a) HPSv2
(b) Aesthetic Score
(c) ImageReward
Figure 14: Illustration of the training cost: training time (GPU hours) against the reward.
D.2.6 Computational Cost of Text-to-Image Generation
Table 6: Comparisons of inference time cost on the text-to-image generation.
# particles 1 2 4 8 16
BoN (s) 3.91 7.10 13.13 24.77 48.47
$\text{SMC}_{\text{base/amot}}$ (s) - 12.20 21.69 41.44 80.61
$\text{SMC}_{\text{grad}}$ (s) 16.19 26.70 47.98 95.70 181.26
To provide a clear picture of the computational cost of the text-to-image experiment, we plot the training time (GPU hours) against reward in Figure˜14. It can be seen that fine-tuning the Meissonic model requires approximately 300 GPU hours using 8 GPUs, which corresponds to roughly 1.5 days of wall-clock time. For the inference cost, we summarise the wall-clock time in Table˜6, where we measure the time required to sample an image from a single prompt using different numbers of particles. The results show that $\text{SMC}_{\text{base}}$ and $\text{SMC}_{\text{amot}}$ methods are more expensive than BoN. This is expected since SMC must evaluate the reward function repeatedly along the sampling trajectory, whereas BoN evaluates it only once. Moreover, $\text{SMC}_{\text{grad}}$ is the most computationally costly, as it requires computing the gradient of the reward at each step.
@@ D.2.6 Computational Cost of Text-to-Image Generation
D.2.7 More Qualitative Results with Generated Samples
@@
  • In Figure˜15, we visualise the generated samples using different methods on ImageReward.
  • In Figure˜16, we visualise the generated samples on HPSv2.
  • In Figure˜17, we visualise the generated samples on Aesthetic Score.
  • In Figure˜18, we visualise the generated samples on ImageReward.
  • In Figure˜19, we demonstrate the generated toxic text using different methods.